Minutes of the Tracker Meeting

July 29, 2004


0) Thermal-Vacuum test preparation

Nicola reported that the inner shield is assembled and is being proof tested.  The inner surface will be painted black at Plyform.  All will be ready at the beginning of September.  He said that the TQCM from Alenia was 20,000 euros for 10 days.

1) Jeff:

      - Plyform EM sidewall coupons with inserts.  Is there still any point in pulling them?  All agreed that we should not bother with pulling them since the design of the sidewalls has changed significantly since then.

      - BR127 primer shipment

2) Sidewall status

Nanda said that so far there is no change to the sidewall spec document.  He will try to close the issue this week.  Plyform has done a test layup.  The result was 1.55 to 1.60 mm thick with the aluminum and very flat.  They have also done the lamination for the coupon tests and will test the coupons next week.

3) Tray panel fabrication status

Two trays will arrive tomorrow from Plyform.  Delivery of all trays is expected next Wednesday.  Friday they will be shake tested.  Over the weekend & Monday the T/V of the panels will be done.  Sandro said that they will do T/V testing of complete trays at G&A for those cases with bubbles or without sandblasting of the tungsten.

The Tower-A schedule will be updated Friday based on current status.

      - Test temperature requirements:

            * can we specify a temperature at which we know our design works?

Sandro said that at room temperature there is no problem while at 55C they have some problems.  Intermediate temperatures have not been investigated in detail.  For the upcoming tray T/V test they will first try a couple of lower temperatures to investigate this issue.

            * what is the maximum tower-tray operational temperature?

4) G&A tray assembly status; report on encapsulation and T/V tests of the first trays

The two trays were subjected to the T/V test at 50C for >12 hours at 200mbar pressure.  There was no change following the test in function or leakage current.  In the next day or so they will do the thermal cycles cold/hot.

Sandro said that there were 3 missing channels from a problem with encapsulation of the first MCM, but they solved the problem after that first layer and the remaining 3 had no problem.

There is one more tray at G&A ready to go and 2 more will be there tomorrow.  Another 2 more will arrive at G&A Tuesday and the rest at the end of next week. 

5) Dave: status of SLAC procurements:

      - I/F hardware

The cones are in hand; hex bushings are out for dycronite; studs are in hand; they completed the inspection of the hex nuts

Arthur is working on the shims: we will use nickel plating; it will be done internally at SLAC

      - Sidewall fasteners

The M2.5 * 100 degree are ready to send

The M4 * 6mm screws are still to arrive

      - Lifting fixture

The first one is ready to ship; INFN needs one more plus a spare (total of 3 in Italy)

      - Inner container

It is assembled; proof loading will be done mid next week.

The cable holding plate is in hand.

      - Outer container

A draft drawing is ready today for checking by Martin and will be send out for review soon, maybe tomorrow.

      - Bending tools

Arthur: early next week he will have the 1st article.

He will make a mockup of the top bend to do tests of the installation; it will be in the shop early next week

      - Top corner brackets

These are in fab and will be ready in another 1.5 weeks

      - Nests

These are at SLAC and need CMM; Jeff needs the specs on the tooling balls.

      - First tower baseplate: it will be in early next week from Humboldt for use on Tower A

      - Some extra heat straps are ready to send to Pisa.

      - 2mm pins are ready to ship (for the bottom tray closeout)

      - SLAC ordered the environmental recorder and digital torque meter

      - The heat strap tie-down strap drawing was released; the strap in Italy needs to be modified accordingly.

6) Robert: EGSE status

      - TEM/PS for thermal/vacuum testing

It is ready from the electronics group and should be tested today by Hiro.

      - EGSE cables

Jerry is working on getting a revised design of the C0 cables shortened to fit on a standard panel and will have them made on a rigid board.  Nanda said that the existing jumpers are long enough to connect such a board into the TEM.  INFN will check on the quantity of jumpers.

7) Tower assembly preparations:

      - Dave: SLAC support plans & travel

Ronaldo questioned the need for so much SLAC support for the alignment procedure.  This will be discussed in the Friday-morning management meeting.

      - Assembly procedure

Emilio is working on it; tomorrow or Monday the next update will be ready for review.

Nicola Saggini is working on the procedure for transportation and handling.

Arthur asked if there are more details on cable installation.  He will send email next week with some suggestions.

      - Marting said that there will be an updated assembly drawing draft today.

      - Alignment procedure

A new draft came out yesterday---it should be nearly ready for release.

      - Date for PRR

Pisa suggested August 25.

      - Electrical test procedure

The 191 document is still in work by Luca.

      - Turn-over fixture status, including proof tests and procedures

Emilio reported that the machine shop is working on it; it will be ready early next week.  He is working on plans for proof tests and will send a draft out tomorrow; Martin said that we need the analysis report for review by John Ku.  Emilio said that the report will include hand calculations.  Martin agree that that was probably sufficient.

      - Date for pre-ship review: not discussed

There was a discussion on how to obtain a measured cg for the tower.  Due to the expense of making a new fixture for a direct measurement, the preference was to weight all of the tower parts, including trays, and calculate the cg.  Martin said that he can put the tray weights into his spread sheet to calculate the cg. 

8) MCM status

      - July 22 status: 194 received at SLAC, 109 reworked

      - 36 out of burn-in yesterday and under test; 27 starting burn-in

      - New pitch adapter status: pitch adapters are at SLAC and Teledyne has the contract to do the qualification work

10) A.O.B.